Compute the payback statistic for Project A if the appropriate cost of capital is 7 percent and the maximum allowable payback period is four years.

Time: 0 1 2 3 4 5
Cash flow: –$1,700 $630 $690 $660 $440 $240

compute payback in years
Answer #1

Payback Period is the time in which initial investment comes back as inward cash flow.

Payback Period = 1(630) + 1(690) + 0.57(380/660)

Payback Period = 2.57 years

Discounted Payback Period,

Year 1, 630/1.07 = $588.78

Year 2, 690/(1.07)2 = $602.67

Year 3, 660/(1.07)3 = $538.75

Year 4, 440/(1.07)4 = $335.67

Discounted Payback Period = 1(588.78) + 1(602.67) + 0.944(508.55/538.75)

Discounted Payback Period = 2.944 years